# QCSE: A Pretrained Quantum Context-Sensitive Word Embedding for Natural Language Processing

### v1

@@ I Introduction
Natural Language Processing (NLP) has rapidly become a cornerstone of modern technology and industry, playing a key role in applications ranging from language translation and text generation, to virtual assistants and data understanding and analysis [1, 2, 3, 4]. These advances have been driven by powerful machine learning models that enable computers to process, understand, reason, and generate human language [5, 6, 7]. Despite this progress, NLP models still face significant challenges, especially when it comes to understanding complex context, meaning, and nuances in language[8, 9]. Classical models are computationally expensive and still face limitations in capturing deep contextual meaning, struggling with certain subtleties in natural language, especially in complex language structures[10, 11, 12, 8, 9, 13, 14, 15, 16, 17, 18]. As the demand for more efficient and accurate NLP solutions grows, researchers are looking for new approaches to push the boundaries. Quantum computing [19, 20, 21, 22] offers a promising frontier in this regard.
@@
With its ability to process information in ways that classical computers cannot, quantum computing opens up new computation possibilities. Quantum machine learning (QML) [23, 24, 25, 26] and QNLP [27, 22, 28] explore how the principles of quantum mechanics — like superposition and entanglement — can be applied to understand language more effectively. In recent years, several pioneering works have laid the foundations for QML in NLP. Following works [29, 30, 25, 31] provide a foundational understanding of QML, explaining how quantum computers could enable new learning algorithms that go beyond classical machine learning. Although QNLP is still in its infancy, its potential to surpass classical embeddings is being actively investigated, potentially leading to breakthroughs in how language is modeled and understood. Pioneering works on QNLP models, such as the quantum language model (QLM) [32, 33] has been explored for a variety of NLP tasks, including question answering and sentiment analysis[34, 35, 36], where superposition and entanglement-based embeddings have shown promising results in capturing dependencies between words in sentences. The rise of QNLP has sparked interest in developing particularly quantum word and sequence embedding models [37, 38, 39, 22, 40].
@@
Quantum word embedding aims to represent words as quantum states that inherently capture their context and meaning for subsequent NLP tasks. A wide range of quantum embedding models has been proposed to handle specific language representation challenges. One such model is the complex-valued embedding model (CVE) [37, 41, 42, 43], which uses quantum-inspired representations to map words onto complex-valued vectors, capturing richer information compared to real-valued embeddings. The CVE model has been applied with promising results to NLP tasks such as sentiment analysis and classification [44, 45, 46]. Similarly, the quantum-inspired pretrained model [47] is designed to integrate quantum principles with classical deep neural networks, offering a hybrid approach to learning word embeddings with pretraining strategies commonly found in NLP. Other recent contributions, which exploit quantum entanglement are word2ket [38] and the quantum language model with entanglement embedding (QLM-EE) [32]. QLM-EE adapts classical pretrained architectures with quantum principles for text understanding. These models mark significant progress in the introduction of quantum-based techniques to NLP and open the door to future innovations in the field.
@@
Despite their potential, current quantum embedding models face limitations that hinder their ability to fully exploit the advantages of quantum computing. One of the most notable limitations is their reliance on pre-trained classical embedding models, such as GloVe, RoBERTa, or word2vec[44, 32, 47] or random initialization, to generate initial word vectors. These learned classical embeddings or vectors are then mapped onto quantum states, rather than training the quantum embedding models directly from the corpus to learn quantum-native and contextual relationships. This dependence could lead to the propagation of the inherent limitations of classical models through the subsequent stages of the quantum process by inheriting the constraints of classical representations.
@@
  • Introduce a quantum embedding model that directly learns the context of words from the corpus, bypassing the need for pre-trained classical word embeddings[44, 32, 47]. Unlike existing models that rely on classical pre-trained word vectors, the proposed quantum model directly learns from the word co-occurrence data in the corpus. This approach enables quantum states to be constructed without dependency on traditional embedding models, offering a fully quantum-native representation of words and their contextual relationships.
@@ II-A Quantum Computing Concepts
To provide a foundational understanding of the key concepts driving quantum computing and QNLP, this section introduces the fundamental principles of bits vs. qubits, superposition, entanglement, and quantum states with $m$ qubits, along with their mathematical formulations [48]. These concepts form the basis for exploring quantum word embeddings and their applications in NLP.
@@
Here, $|0\rangle$ and $|1\rangle$ are the basis states of the qubit.
@@
Entanglement[49, 50] is a quantum phenomenon in which two or more quantum systems become correlated in such a way that the state of one cannot be described independently of the others. The most simple relevant entangled states are the Bell states. One of these Bell states can be written as
@@ II-B Quantum Natural Language Processing
Natural Language Processing involves two key stages: word embedding and model training. Given a vocabulary $V$, words are mapped to $d$-dimensional vectors via an embedding function $f:V\rightarrow\mathbb{R}^{d}$. Popular methods include word2vec [51] and GloVe [52] for maximizing the log-likelihood of context words $w_{c}$ given center word $w$:
@@
These embeddings $x_{1},...,x_{T}$} are then processed by sequence models such as RNNs/LSTMs [53, 54, 55] and transformers [56, 57] for learning patterns over sequences of words. In a recurrent model like LSTM, the hidden states $h_{t}$ are computed as a function of the previous hidden states and the current input: $h_{t}=f(h_{t-1},x_{t})$. $h_{t}=f(h_{t-1},x_{t})$. The training minimizes for example a cross-entropy loss $\mathcal{L}_{\text{train}}=-\sum_{i=1}^{N}y_{i}\log\hat{y}_{i}$, where $y_{i}$ is the true label and $\hat{y}_{i}$ is the predicted probability for sample $i$.
@@
QNLP leverages the principles of quantum computing to enhance traditional NLP tasks. A key concept in QNLP is the use of quantum word embedding, where classical data (words or sentences) are embedded into quantum states. Instead of representing words classically in a high-dimensional vector space, each word is encoded in a quantum state $|\psi\rangle$ in a Hilbert space. Mathematically, this can be seen as mapping each word $w$ in the vocabulary $V$ to a quantum state $|\psi_{w}\rangle$, where the amplitude or angle of each qubit encodes information about the context of the word. For instance, using amplitude encoding, a vector $\mathbf{v}\in\mathbb{R}^{d}$ representing a word can be normalized and encoded into a quantum state $|\psi\rangle=\sum_{i=1}^{d}v_{i}|i\rangle$. Quantum word embeddings could leverage the superposition and entanglement properties of quantum systems to capture richer contextual relationships between words than classical embeddings.
@@
Training QNLP models typically involves a QML framework such as the variational quantum circuit (VQC)[58, 29, 59], where the circuit parameters are optimized to minimize a loss function. Let $|\psi_{w}\rangle$ be the initial quantum state representing a word or sentence. The VQC transforms this state into another state through a series of unitary operations, $|\psi_{\text{output}}\rangle=U(\theta)|\psi_{w}\rangle$, where $U(\theta)$ is a parameterized unitary matrix dependent on the parameters $\theta$. These parameters are updated via classical optimization algorithms, such as gradient descent, minimizing a loss function $\mathcal{L}(\theta)$ computed from quantum measurements. In QNLP, such training often involves measuring the expectation values of quantum operators to compute probabilities for different outcomes, similar to classical model training. In the case of the quantum embedding model, the final model aims at maximizing the likelihood $P(|w_{\text{context}}\rangle||w_{\text{center}\rangle},\theta)$, where $|w_{\text{center}}\rangle$ and $|w_{\text{context}}\rangle$ are quantum states of the word to embed and its context words in a sentence, respectively, similar to the classical word embedding prediction tasks. This enables the model to learn contextual word embeddings in a quantum-enhanced space. Quantum encoding strategies are crucial for effectively transforming text into processable quantum states, offering potential advantages over classical methods in contextual representation.
@@ III-A Model overview
The aim is to map each word to a quantum state that captures its semantic meaning in the context of a given task. Suppose that we have a set of words $W=\{w_{1},w_{2},\dots,w_{n}\}$, and a vocabulary of size $|V|$. In classical word embeddings, each word is mapped to a high-dimensional vector $\mathbf{w}_{i}\in\mathbb{R}^{d}$, where $d$ is the embedding dimension. For quantum embeddings, the quantum model instead maps words to quantum states $|w_{i}\rangle\in\mathcal{H}_{m}$, where $\mathcal{H}_{m}$ is a Hilbert space of dimension $m$, and the quantum embedding process is represented by a unitary transformation that incorporates the word context. The process of learning quantum embeddings involves applying a VQC, typically referred to as an ansatz, to capture the relationships between words in the context of a given task. The ansatz parameters are updated iteratively using optimization methods, such as gradient-based optimization, to minimize a task-specific loss function, such as the cross-entropy. For a quantum sequence embedding model, words in a sentence are encoded as sequences of quantum states, and entanglement operations can capture the dependencies between words in a sequence. We propose a context matrix concept (see Section IV) that transforms the surrounding terms of a word into a structured matrix that captures identifiers, positions, distances, and semantic–syntactic relations to construct the context-sensitive quantum embedding.
Figure 1: Quantum Embedding Model for Word/Sequence Embedding. The model combines: (1) a quantum context-encoding circuit, and (2) a parameterized quantum circuit trained to generate context-sensitive embeddings. Step 1: Define the context {natural, language, is, awesome} and identify the center word (here, processing). The context is transformed into a context matrix. Step 2: Encode the matrix using the context encoding circuit. Context matrices are reshaped for quantum encoding (no resizing required when columns match qubit count). The flexible design adapts to diverse NLP tasks. Step 3: The model predicts the quantum embedding of the center word. Steps 4–6: Compute the loss against the true embedding of the center word.
@@ III-B1 Quantum Context Embedding Circuit
Subsequently, each vector $\mathbf{\tilde{c}}_{l}$ in the context matrix encodes the rotation angles for the quantum gates, as the $l$-th layer of the encoding circuit. Specifically, for each qubit $q\in\{1,\cdots,m\}$, we apply two rotation gates $RX_{q}(\mathbf{\tilde{c}}_{l}^{(q_{1})})$ and $RZ_{q}(\mathbf{\tilde{c}}_{l}^{(q_{2})})$ around the $x$ and $z$-axis of the Bloch sphere, as follows:
@@
$=\begin{pmatrix}\cos(\mathbf{\tilde{c}}_{l}^{(q_{1})}/2)&-i\sin(\mathbf{\tilde{c}}_{l}^{(q_{1})}/2)\\ -i\sin(\mathbf{\tilde{c}}_{l}^{(q_{1})}/2)&\cos(\mathbf{\tilde{c}}_{l}^{(q_{1})}/2)\end{pmatrix}$
@@
$=\begin{pmatrix}\exp(-i\mathbf{\tilde{c}}_{l}^{(q_{2})}/2)&0\\ 0&\exp(i\mathbf{\tilde{c}}_{l}^{(q_{2})}/2)\end{pmatrix}$
where $\mathbf{\tilde{c}}_{l}^{(q_{1})}\text{ and }\mathbf{\tilde{c}}_{l}^{(q_{2})}$ represent the two context parameters for qubit $q$ in each encoding layer. This encoding operation is repeated for each vector $\mathbf{\tilde{c}}_{l}$ in the context matrix, ensuring that the input context is fully encoded into the quantum state.
The full quantum encoding operation for each layer $\mathbf{\tilde{c}}_{l}$ of the context matrix can be described as:
$\lvert\psi_{\text{enc}}\rangle=\prod_{q=1}^{m}\text{$RZ_{q}$}(\mathbf{\tilde{c}}_{l}^{(q_{1})})\text{$RX_{q}$}(\mathbf{\tilde{c}}_{l}^{(q_{2})})\lvert\psi_{0}\rangle,$ (13)
@@
$\lvert\psi_{\text{ent}}\rangle=\prod_{q=1}^{m-1}\text{CNOT}(q,q+1)\lvert\psi_{\text{enc}}\rangle.$ (14)
@@
For illustration purposes, we have added an example in Figure 1. The example illustrates the process of transitioning from a word, processing in this case, and defining its context (natural, language, is, awesome) within the corpus using the bag-of-word approach [60] to define the context of the word (another process can be used, for example, using a character-level process). Then, the context, based on the indexes, positions, frequencies, and distances of words, is transformed into a matrix which is then encoded in a quantum state. This context-sensitive encoding passes through the ansatz for predicting the quantum word embedding, within its context.

### v2

@@ I Introduction
Natural Language Processing (NLP) has rapidly become a cornerstone of modern technology and industry, playing a key role in applications ranging from language translation and text generation, to virtual assistants and data understanding and analysis [uddin2024review, pais2022nlp, lane2025natural, tyagi2023demystifying]. These advances have been driven by powerful machine learning models that enable computers to process, understand, reason, and generate human language [kalyan2021ammus, treviso2023efficient, qin2024large]. Despite this progress, NLP models still face significant challenges, especially when it comes to understanding complex context, meaning, and nuances in language[khan2023exploring, hadi2023large]. Classical models are computationally expensive and still face limitations in capturing deep contextual meaning, struggling with certain subtleties in natural language, especially in complex language structures[mars2022word, gao2021limitations, wang2020survey, khan2023exploring, hadi2023large, lin2023artificial, koubaa2023exploring, li2020openai, a3, a4, a5]. As the demand for more efficient and accurate NLP solutions grows, researchers are looking for new approaches to push the boundaries. Quantum computing [guarasci2022quantum, widdows2024quantum, abbaszade2021application, coecke2020foundations] offers a promising frontier in this regard.
@@
With its ability to process information in ways that classical computers cannot, quantum computing opens up new computation possibilities. Quantum machine learning (QML) [zhang2020recent, narwane2025quantum, wang2024comprehensive, olaitan2025quantum] and QNLP [varmantchaonala2024quantum, coecke2020foundations, widdows2024near] explore how the principles of quantum mechanics — like superposition and entanglement — can be applied to understand language more effectively. In recent years, several pioneering works have laid the foundations for QML in NLP. Following works [schuld2021machine, schuld2015introduction, wang2024comprehensive, peral2024systematic] provide a foundational understanding of QML, explaining how quantum computers could enable new learning algorithms that go beyond classical machine learning. Although QNLP is still in its infancy, its potential to surpass classical embeddings is being actively investigated, potentially leading to breakthroughs in how language is modeled and understood. Pioneering works on QNLP models, such as the quantum language model (QLM) [chen2021quantum, di2022dawn] has been explored for a variety of NLP tasks, including question answering and sentiment analysis[zhang2018end, zhang2018quantum, zhang2020quantum], where superposition and entanglement-based embeddings have shown promising results in capturing dependencies between words in sentences. The rise of QNLP has sparked interest in developing particularly quantum word and sequence embedding models [li2018quantum, panahi2019word2ket, nguyen2024quantum, coecke2020foundations, karamlou2022quantum].
@@
Quantum word embedding aims to represent words as quantum states that inherently capture their context and meaning for subsequent NLP tasks. A wide range of quantum embedding models has been proposed to handle specific language representation challenges. One such model is the complex-valued embedding model (CVE) [li2018quantum, harvey2024learning, zhang2022complex, jaiswal2018quantum], which uses quantum-inspired representations to map words onto complex-valued vectors, capturing richer information compared to real-valued embeddings. The CVE model has been applied with promising results to NLP tasks such as sentiment analysis and classification [chu2024effective, zhao2022quantum, shi2021two]. Similarly, the quantum-inspired pretrained model [shi2024pretrained] is designed to integrate quantum principles with classical deep neural networks, offering a hybrid approach to learning word embeddings with pretraining strategies commonly found in NLP. Other recent contributions, which exploit quantum entanglement are word2ket [panahi2019word2ket] and the quantum language model with entanglement embedding (QLM-EE) [chen2021quantum]. QLM-EE adapts classical pretrained architectures with quantum principles for text understanding. These models mark significant progress in the introduction of quantum-based techniques to NLP and open the door to future innovations in the field.
@@
Despite their potential, current quantum embedding models face limitations that hinder their ability to fully exploit the advantages of quantum computing. One of the most notable limitations is their reliance on pre-trained classical embedding models, such as GloVe, RoBERTa, or word2vec[chu2024effective, chen2021quantum, shi2024pretrained] or random initialization, to generate initial word vectors. These learned classical embeddings or vectors are then mapped onto quantum states, rather than training the quantum embedding models directly from the corpus to learn quantum-native and contextual relationships. This dependence could lead to the propagation of the inherent limitations of classical models through the subsequent stages of the quantum process by inheriting the constraints of classical representations.
@@
  • Introduce a quantum embedding model that directly learns the context of words from the corpus, bypassing the need for pre-trained classical word embeddings[chu2024effective, chen2021quantum, shi2024pretrained]. Unlike existing models that rely on classical pre-trained word vectors, the proposed quantum model directly learns from the word co-occurrence data in the corpus. This approach enables quantum states to be constructed without dependency on traditional embedding models, offering a fully quantum-native representation of words and their contextual relationships.
@@ II-A Quantum Computing Concepts
To provide a foundational understanding of the key concepts driving quantum computing and QNLP, this section introduces the fundamental principles of bits vs. qubits, superposition, entanglement, and quantum states with $m$ qubits, along with their mathematical formulations [nielsen2010quantum]. These concepts form the basis for exploring quantum word embeddings and their applications in NLP.
@@
Here, $|0\rangle$ and $|1\rangle$ are the computational basis of the qubit.
@@
Entanglement[horodecki2009quantum, guhne2009entanglement] is a quantum phenomenon in which two or more quantum systems become correlated in such a way that the state of one cannot be described independently of the others.
The most simple relevant entangled states are the Bell states. One of these Bell states can be written as
@@ II-B Quantum Natural Language Processing
Natural Language Processing involves two key stages: word embedding and model training. Given a vocabulary $V$, words are mapped to $d$-dimensional vectors via an embedding function $f:V\rightarrow\mathbb{R}^{d}$. Popular methods include word2vec [mikolov2013distributed] and GloVe [pennington2014glove] for maximizing the log-likelihood of context words $w_{c}$ given center word $w$:
@@
These embeddings $x_{1},...,x_{T}$} are then processed by sequential models such as RNNs/LSTMs [tarwani2017survey, xiao2020research, yao2018improved] and transformers [gillioz2020overview, patwardhan2023transformers] for learning patterns over sequences of words. In a recurrent model like LSTM, the hidden states $h_{t}$ are computed as a function of the previous hidden states and the current input: $h_{t}=f(h_{t-1},x_{t})$. The training minimizes for example a cross-entropy loss $\mathcal{L}_{\text{train}}=-\sum_{i=1}^{N}y_{i}\log\hat{y}_{i}$, where $y_{i}$ is the true label and $\hat{y}_{i}$ is the predicted probability for sample $i$.
@@
QNLP leverages the principles of quantum computing to enhance traditional NLP tasks. A key concept in QNLP is the use of quantum word embedding, where classical data (words or sentences) are embedded into quantum states. Instead of representing words as classical high-dimensional vectors, each word is encoded as a quantum state $|\psi\rangle$ in a Hilbert space. In this approach, every word $w$ in the vocabulary $V$ is mapped to a quantum state $|\psi_{w}\rangle$, where the probability amplitudes of the qubits store information about the word. For example, with amplitude encoding, a classical word vector $\mathbf{v}\in\mathbb{R}^{d}$ can be normalized and stored in a quantum state as $|\psi\rangle=\sum_{i=1}^{d}v_{i}|i\rangle$. Quantum word embeddings could leverage the superposition and entanglement properties of quantum systems to capture richer contextual relationships between words than classical embeddings.
@@
Training QNLP models typically involves a QML framework such as the variational quantum circuit (VQC)[cerezo2021variational, schuld2021machine, havlivcek2019supervised], where the circuit parameters are optimized to minimize a loss function. Let $|\psi_{w}\rangle$ be the initial quantum state representing a word or sentence. The VQC transforms this state into another state through a series of unitary operations, $|\psi_{\text{output}}\rangle=U(\theta)|\psi_{w}\rangle$, where $U(\theta)$ is a parameterized unitary matrix dependent on the parameters $\theta$. These parameters are updated via classical optimization algorithms, such as gradient descent, minimizing a loss function $\mathcal{L}(\theta)$ computed from quantum measurements. In QNLP, such training often involves measuring the expectation values of quantum operators to compute probabilities for different outcomes, similar to classical model training. In the case of the quantum embedding model, the final model aims at maximizing the likelihood $P(|w_{\text{context}}\rangle||w_{\text{center}\rangle},\theta)$, where $|w_{\text{center}}\rangle$ and $|w_{\text{context}}\rangle$ are quantum states of the word to embed and its context words in a sentence, respectively, similar to the classical word embedding prediction tasks. This enables the model to learn contextual word embeddings in a quantum-enhanced space. Quantum encoding strategies are crucial for effectively transforming text into processable quantum states, offering potential advantages over classical methods in contextual representation.
@@ III-A Model overview
The aim is to map each word to a quantum state that captures its semantic meaning in the context of a given task. Suppose that we have a set of words $W=\{w_{1},w_{2},\dots,w_{n}\}$, and a vocabulary of size $|V|$. In classical word embeddings, each word is mapped to a high-dimensional vector $\mathbf{w}_{i}\in\mathbb{R}^{d}$, where $d$ is the embedding dimension. For quantum embeddings, the quantum model instead maps words to quantum states $|w_{i}\rangle\in\mathcal{H}$, where $\mathcal{H}$ is a Hilbert space, and the quantum embedding process is represented by a unitary transformation that incorporates the word context. The process of learning quantum embeddings involves applying a VQC, typically referred to as an ansatz, to capture the relationships between words in the context of a given task. The ansatz parameters are updated iteratively using optimization methods, such as gradient-based optimization, to minimize a task-specific loss function, such as the cross-entropy. We propose a context matrix concept (see Section IV) that transforms the surrounding terms of a word into a structured matrix that captures identifiers, positions, distances, and semantic–syntactic relations to construct the context-sensitive quantum embedding.
Figure 1: QCSE architecture and training pipeline. The model combines: (1) a quantum context-encoding circuit, and (2) a parameterized quantum circuit trained to generate context-sensitive embeddings. Step 1: Define the context {natural, language, is, awesome} and identify the center word (here, processing). The context is transformed into a context matrix. Step 2: Encode the matrix using the context encoding circuit. Context matrices are reshaped for quantum encoding (no resizing required when columns match qubit count). The flexible design adapts to diverse embedding approaches. Step 3: The model predicts the quantum embedding of the center word. Steps 4–6: Compute the loss against the true embedding of the center word and optimize ansatz parameters.
@@ III-B1 Quantum Context Embedding Circuit
Subsequently, each vector $\mathbf{\tilde{c}}_{l}$ in the context matrix encodes the rotation angles for the quantum gates, as the $l$-th layer of the encoding circuit. Specifically, for each qubit $q\in\{1,\cdots,m\}$, we apply two rotation gates $RX_{q}(\mathbf{\tilde{c}}_{l}^{(q,1)})$ and $RZ_{q}(\mathbf{\tilde{c}}_{l}^{(q,2)})$ around the $x$ and $z$-axis of the Bloch sphere, as follows:
@@
$=\begin{pmatrix}\cos(\mathbf{\tilde{c}}_{l}^{(q,1)}/2)&-\text{i}\sin(\mathbf{\tilde{c}}_{l}^{(q,1)}/2)\\ -\text{i}\sin(\mathbf{\tilde{c}}_{l}^{(q,1)}/2)&\cos(\mathbf{\tilde{c}}_{l}^{(q,1)}/2)\end{pmatrix}$
@@
$=\begin{pmatrix}\exp(-\text{i}\mathbf{\tilde{c}}_{l}^{(q,2)}/2)&0\\ 0&\exp(\text{i}\mathbf{\tilde{c}}_{l}^{(q,2)}/2)\end{pmatrix}$
where $\mathbf{\tilde{c}}_{l}^{(q,1)}\text{ and }\mathbf{\tilde{c}}_{l}^{(q,2)}$ represent the two context parameters for qubit $q$ in each encoding layer. This encoding operation is repeated for each vector $\mathbf{\tilde{c}}_{l}$ in the context matrix, ensuring that the input context is fully encoded into the quantum state.
The complete rotation operations for each layer $\mathbf{\tilde{c}}_{l}$ of the context matrix can be described as:
$\lvert\psi_{\text{enc}}\rangle=\left(\bigotimes_{q=1}^{m}RZ_{q}(\tilde{c}_{l}^{(q,1)})RX_{q}(\tilde{c}_{l}^{(q,2)})\right)\lvert\psi_{0}\rangle$ (13)
@@
$\lvert\psi_{\text{ent}}\rangle=\text{CNOT}(m-1,m)\cdots\text{CNOT}(2,3)\cdot\text{CNOT}(1,2)\lvert\psi_{\text{enc}}\rangle$ (14)
@@
For illustration purposes, we have added an example in Figure 1. The example illustrates the process of transitioning from a word, processing in this case, and defining its context (natural, language, is, awesome) within the corpus using the bag-of-word approach [sivakumar2020review] to define the context of the word (another process can be used, for example, using a character-level process). Then, the context, based on the indexes, positions, frequencies, and distances of words, is transformed into a matrix which is then encoded in a quantum state. This context-sensitive encoding passes through the ansatz for predicting the quantum word embedding, within its context.